# Dormant BH candidates from Gaia DR3 summary diagnostics

### v1

@@ 2.2 Spectroscopy
As a spectroscopic variability indicator, we forward-model the formal “uncertainty” on the median RV reported in the Gaia DR3 catalogue. For a star in a Keplerian binary, the RV varies over time according to
@@
where $\nu$ is the true anomaly, $e$ is the eccentricity, and $v_{z}$ is a constant velocity offset. The RV semi-amplitude $K_{*}$ of the primary star induced by the companion is a function of the companion mass, orbital period, eccentricity, and inclination;
@@
The RV amplitude increases with larger companion masses and smaller separations of the stars.
We use gaiamock to compute epoch-level RVs by sampling Equation 2 at realistic Gaia-like observing times. Because the precise spectroscopic scanning law is not public, we approximate the RV observing epochs using a random subset of rv_nb_transits times from the observing dates given by the astrometric scanning law. While not exact, this approximation appears sufficient for our purposes: with a median of $\sim$18 transits per star in DR3 (Katz et al., 2023), binaries with periods shorter than or comparable to the Gaia baseline are typically sampled well enough that the specific observing times have little effect on the variability statistics.
@@
where $N$ is the number of transits (rv_nb_transits), and the additive term accounts for the spectroscopic noise floor from wavelength calibration uncertainties (Katz et al., 2023). We compute the standard deviation from the simulated RVs and apply this transformation to predict the expected radial_velocity_error, which we refer to as the RV uncertainty $\sigma_{\mathrm{RV}}$.
@@
To validate our spectroscopic variability model and empirically calibrate its uncertainties, we tested it on known spectroscopic binaries from the Gaia NSS catalogue, selecting systems with nss_solution_type = SB1. As long as our predictions include only the RV variability induced by binary motion, neglecting measurement errors and additional noise sources, they will underestimate the reported RV uncertainties. To account for this, we applied a uniform correction factor derived from the observed biases in the validation sample and scaled the predicted RV uncertainties by a factor of 1.07. Fig. 2 compares our predicted versus the catalogued RV uncertainties. The tight linear correlation confirms that the model captures the key trends in this metric. As in the astrometric case, the scatter in the logarithmic residuals is approximately constant. We therefore adopt fractional uncertainties of $\epsilon_{\mathrm{RV}}=0.15$.
@@ 2.3 Photometry
To model photometric variability, we adopt an analytical prescription for the amplitude of ellipsoidal variations, that is, the brightness changes arising from tidal distortion of a star in a close binary system (Morris & Naftilan, 1993). These effects are strongest at short orbital periods and diminish rapidly when the binary separation is much larger than the radii of the component stars. Our aim is not to capture the variability in detail but to estimate its amplitude to order-of-magnitude precision, sufficient to exclude short-period binaries from our candidate sample.
Morris & Naftilan (1993) derive an analytical model for ellipsoidal modulation in circular orbits, expressed as a power series in $R_{*}/a$, the ratio of the primary star’s radius to the system’s semimajor axis. We use the dominant second-harmonic term from their equation (1), incorporating the correction factor from Gomel et al. (2021). Under the assumption that the primary star dominates the flux, the semi-amplitude of the photometric variability $A_{\mathrm{phot}}$ can then be expressed as a function of the binary mass ratio $q$, inclination angle, and Roche-lobe filling factor of the primary (which depends on $R_{*}$, $M_{*}$, $q$, and $a$):
@@
The coefficients $\alpha_{2}$ and $\beta_{2}$, along with the correction factor $c_{2}$ (see Equations (2) and (11) in Gomel et al., 2021), account for limb and gravity darkening, and depend on the stellar effective temperature and surface gravity. We adopt approximate coefficient values consistent with the Gaia BH2 analysis (El-Badry et al., 2023a), which is adequate given our focus on detached binaries containing cool stars with convective envelopes.
@@ 3 Binary parameter inference
In Section 2, we showed that for known binary parameters, we can reliably predict the Gaia observables, in particular the renormalised unit weight error (ruwe) and the RV uncertainty ($\sigma_{\mathrm{RV}}$). We now address the inverse problem: How well can one constrain binary parameters from these summary diagnostics using a forward-modelling Bayesian framework.
We adopt a Markov Chain Monte Carlo (MCMC) approach, using the emcee sampler (Foreman-Mackey et al., 2013) to sample the posterior distribution of binary parameters given observed Gaia summary statistics. Specifically, we infer the posterior probability distribution of the parameter vector
$\@vec{\theta}=\left(P,M_{\text{CO}},\cos i,e,\Omega,\omega,T_{\text{p}}\right),$
@@
$\@vec{s_{*}}=\left(M_{*},T_{\text{eff,}*},R_{*},G,\varpi,\alpha,\delta\right).$
The seven orbital parameters cannot be uniquely constrained by the available observables alone. However, we assess whether meaningful constraints can still be obtained on the most relevant parameters by marginalising over the other ones. Our primary parameters of interest are the orbital period ($P$) and the unseen companion mass ($M_{\text{CO}}$). The remaining orbital parameters – inclination ($i$), eccentricity ($e$), longitude of the ascending node ($\Omega$), argument of periastron ($\omega$), and time of periastron passage ($T_{\text{p}}$) – are treated as nuisance parameters required for forward modelling but marginalised in the final inference.
@@
where ‘obs’ denotes the catalogued summary statistics and ‘pred’ the model predictions for a given parameter set $\@vec{\theta}$. The likelihood measures discrepancies in units of scatter, given by $\epsilon_{\texttt{ruwe}}$ and $\epsilon_{\sigma_{\text{RV}}}$, and penalises them in log-space. The use of logarithmic quantities is empirically motivated; forward modelling of NSS systems shows that the combined scatter – from intrinsic variation and model uncertainty – is constant in log-space and can be approximated by a Gaussian distribution (see Fig. 2). Expressing the likelihood in terms of $\ln\mathtt{ruwe}$ and $\ln\sigma_{\mathrm{RV}}$ also ensures symmetric treatment of relative over- and under-predictions.
@@
We adopt log-uniform priors for period and companion mass, and use uniform priors for all other parameters over physically motivated ranges: $\log P/\mathrm{d}\in[1,4]$, $\log M_{\text{CO}}/M_{\odot}\in[\log(0.05),\log(30)]$, $\cos i\in(0,1]$, $e\in[0,0.95]$, $\Omega,\omega\in[0,2\pi)$, and $T_{p}/\text{d}\in[0,P]$. The flux ratio is fixed to zero, which is appropriate for dark companions. For luminous companions, a non-zero flux ratio would reduce the observed ruwe, leading to underestimated periods and masses. This bias is acceptable since we target dark, massive companions, and for luminous giants, the flux contribution from stellar companions is typically negligible.
@@
We sample posterior distributions using 32 walkers and 2,500 steps with a 500-step burn-in phase. Convergence assessment is challenging due to the underconstrained problem and high shot noise (e.g. from stochastic Gaia RV observation date sampling). We evaluated convergence empirically on a subset of stars by verifying that: (a) randomly initialised walkers converge to consistent posterior distributions, (b) marginal distributions of $P$ and $M_{\text{CO}}$ remain stable between the second and final thirds of the chain, and (c) increasing chain length by a factor of 10 does not significantly alter results. The inference is computationally efficient, requiring approximately 10 seconds per star on a single CPU core. We report posterior-weighted mean values of $P$ and $M_{\text{CO}}$, marginalised over nuisance parameters.
@@
Figure 3 illustrates the results of this MCMC fitting procedure for two test binary systems: a stellar binary (left panels) and a system containing a compact object (right panels), specifically Gaia BH 2 (El-Badry et al., 2023a). The upper panels display the marginalised posterior probability distribution function (PDF) in the plane of orbital period versus companion mass. The orange star denotes the true input parameters of each system, while the blue star marks the inferred best-fit parameters, defined as the posterior-weighted mean. Dashed ellipses show the 1-$\sigma$ and 2-$\sigma$ confidence contours, derived from the covariance matrix of the posterior. Above each panel, we show the marginalised 1-D posterior PDF of the companion mass, normalised so that the sum of all posterior weights is unity. The bottom panels show the predicted observables ruwe and $\sigma_{\text{RV}}$ for each MCMC sample, colour coded by the corresponding companion mass. The orange star shows the observed values (from the Gaia source catalogue).
@@
The first system (Gaia DR3 4609711364265062016) appears in the NSS catalogue as an AstroSpectroSB1 binary with an orbital period of 141 d. The Gaia binary mass catalogue assigns it a primary mass of $\sim 1.1\,M_{\odot}$ and a companion mass of $\sim 0.35\,M_{\odot}$. The observed values of ruwe and RV uncertainty are modest ($\texttt{ruwe}=1.81$, $\sigma_{\text{RV}}=2.10$ km s-1), consistent with a luminous stellar companion. Our inference correctly recovers a low-mass companion in a short-period orbit ($P_{\text{fit}}=110^{+90}_{-50}\,\text{d},\,\,M_{\text{CO, fit}}=0.5^{+0.3}_{-0.2}\,\text{M}_{\odot}$). Small mismatches between the true and inferred values and large uncertainties are expected, given the degeneracies in orbital orientation and eccentricity.
@@
The second example is Gaia BH 2 (Gaia DR3 5870569352746779008), a giant star orbited by a $\sim 9\,\text{M}_{\odot}$ BH in a long-period eccentric orbit ($P\approx 1277$ d, $e\approx 0.53$; El-Badry et al., 2023a). It has a high ruwe value of $9.22$, with moderate RV uncertainty ($\sigma_{\text{RV}}=2.08$ km s-1). Matching both observables requires a high companion mass, and our inference yields a posterior peak at $P_{\text{fit}}=1300^{+1100}_{-600}\,\text{d}$ and $M_{\text{CO, fit}}=9^{+7}_{-4}\,\text{M}_{\odot}$, consistent with the true values and a dark remnant companion. For systems with orbital periods longer than the DR3 baseline, there is a degeneracy between period and companion mass; a longer-period, more massive companion can produce identical observables as a shorter-period, less massive one due to limited orbital coverage. We cap the prior at $30\,\text{M}_{\odot}$, which constrains the posterior to a physically motivated BH mass regime, but also truncates the distribution. For candidates like BH 2 that lie near this boundary, the inferred companion masses should therefore be interpreted as lower limits under the assumed log-uniform prior.
@@ 4 A catalogue of RGB+BH candidates
We are now in a position to apply this forward-modelling technique to a sample of evolved stars from Gaia DR3. Giant stars offer several advantages in searches for compact companions: (1) with their high luminosities they are likely to outshine their companions, resulting in flux ratios near zero and photocentre motions dominated by the primary; and (2) their large current or past radii preclude short-period orbits, reducing contamination from hierarchical triples, a major source of false positives in main-sequence binary searches (Andrew et al., 2022).
@@ 4.1 Sample selection and filtering
These cuts reduced the sample to 37,421 stars. The RV amplitude threshold roughly corresponds to $\sigma_{\mathrm{RV}}\gtrsim 2$ km s-1.These thresholds are motivated by two considerations: (a) In broader tests, systems with low RV scatter and low ruwe – occupying the lower-left region of the bottom panels in Fig. 3 – were almost exclusively binaries with low-mass companions. Due to the rarity of BHs and the high contamination in this region, we focussed on areas of parameter space where BHs exhibit stronger and more distinct dynamical signatures. (b) Although BHs in wide or low-inclination orbits may be excluded by these cuts, we found that the impact on overall completeness is small for the period range of interest (see Sect. 4.5).
To further refine the sample, we cross-matched with the stellar parameter catalogue from Zhang et al. (2023), which provides $T_{\mathrm{eff}}$, $\log g$, and [Fe/H] estimates from BP/RP spectra for 220 million stars. We retained stars with $T_{\mathrm{eff}}<6000$ K to remove reddened main-sequence stars, and required reliable parameter estimates (quality_flags ¡ 8). This removed very red supergiants and reduced the sample to 21,028 evolved stars.
@@
Figure 4: CMD of the RGB candidate sample (2D histogram), overlaid on all stars within 100 pc from Gaia DR3 (grey points). Highlighted are confirmed BH binaries Gaia BH 2 and BH 3, and the candidate LAMOST BH. Neutron star binaries (triangles) and Gaia BH 1 are shown for reference, but lie outside our RGB selection.

### v2

@@ 2.2 Spectroscopy
As a spectroscopic variability indicator, we forward model the formal uncertainty on the median RV reported in the Gaia DR3 catalogue. For a star in a Keplerian binary, the RV varies over time according to
@@
where $\nu$ is the true anomaly, $e$ is the eccentricity, and $v_{z}$ is a constant velocity offset. The RV semi-amplitude $K_{*}$ of the primary star induced by the companion is a function of the companion mass, orbital period, eccentricity, and inclination:
@@
The RV amplitude increases with higher companion masses and smaller separations of the stars.
We use gaiamock to compute epoch-level RVs by sampling Eq. 2 at realistic Gaia-like observation times. Because the precise spectroscopic scanning law is not public, we approximate the RV observing epochs using a random subset of rv_nb_transits times from the observing dates given by the astrometric scanning law. While not exact, this approximation appears sufficient for our purposes: with a median of $\sim$18 transits per star in DR3 (Katz et al., 2023), binaries with periods shorter than or comparable to the Gaia baseline are typically sampled well enough that the specific observing times have little effect on the variability statistics.
@@
where $N$ is the number of transits (rv_nb_transits), and the additive term accounts for the spectroscopic noise floor from wavelength calibration uncertainties (Katz et al., 2023). We compute the standard deviation from the simulated RVs and apply this transformation to predict the expected radial_velocity_error, which we refer to as RV uncertainty $\sigma_{\mathrm{RV}}$.
@@
To validate our spectroscopic variability model and empirically calibrate its uncertainties, we tested it on known spectroscopic binaries from the Gaia NSS catalogue, selecting systems with nss_solution_type = SB1. Because our predictions included only the RV variability induced by binary motion, neglecting measurement errors and additional noise sources, they would underestimate the reported RV uncertainties. To account for this, we applied a uniform correction factor derived from the observed biases in the validation sample and scaled the predicted RV uncertainties by a factor of 1.07. Figure 2 compares our predicted vs the catalogued RV uncertainties. The tight linear correlation confirms that the model captures the key trends in this metric. As in the astrometric case, the scatter in the logarithmic residuals is approximately constant. We therefore adopt fractional uncertainties of $\epsilon_{\mathrm{RV}}=0.15$.
@@ 2.3 Photometry
To model photometric variability, we adopt an analytical prescription for the amplitude of ellipsoidal variations, that is, the brightness changes arising from tidal distortion of a star in a close binary system (Morris and Naftilan, 1993). These effects are strongest at short orbital periods and diminish rapidly when the binary separation is much larger than the radii of the component stars. Our aim is not to capture the variability in detail but to estimate its amplitude to order-of-magnitude precision, sufficient to exclude short-period binaries from our candidate sample.
Morris and Naftilan (1993) derive an analytical model for ellipsoidal modulation in circular orbits, expressed as a power series in $R_{*}/a$, the ratio of the primary star’s radius to the system’s semimajor axis. We use the dominant second-harmonic term from their Eq. (1), incorporating the correction factor from Gomel et al. (2021). Under the assumption that the primary star dominates the flux, the semi-amplitude of the photometric variability $A_{\mathrm{phot}}$ can then be expressed as a function of the binary mass ratio $q$, inclination angle, and Roche-lobe filling factor of the primary (which depends on $R_{*}$, $M_{*}$, $q$, and $a$):
@@
The coefficients $\alpha_{2}$ and $\beta_{2}$, along with the correction factor $c_{2}$ (see Eqs. 2 and 11 in Gomel et al., 2021), account for limb and gravity darkening, and depend on the stellar effective temperature and surface gravity. We adopted approximate coefficient values consistent with a star with properties of the companion to Gaia BH 2 (Claret and Bloemen, 2011; El-Badry et al., 2023a), which is adequate given our focus on detached binaries containing cool stars with convective envelopes.
@@ 3 Binary parameter inference
In Sect. 2 we showed that for known binary parameters, we can reliably predict Gaia observables, in particular the re-normalised unit weight error (ruwe) and the RV uncertainty ($\sigma_{\mathrm{RV}}$). We now address the inverse problem: How well can one constrain binary parameters from these summary diagnostics using a forward-modelling Bayesian framework.
We adopted a Markov chain Monte Carlo (MCMC) approach, using the emcee sampler (Foreman-Mackey et al., 2013), to sample the posterior distribution of binary parameters given the observed Gaia summary statistics. Specifically, we inferred the posterior probability distribution of the parameter vector
$\@vec{\theta}=\left(P,M_{\text{CO}},\cos i,e,\Omega,\omega,T_{\text{p}},\varpi\right),$
@@
$\@vec{s_{*}}=\left(M_{*},T_{\text{eff,}*},R_{*},G,\alpha,\delta\right).$
The orbital parameters cannot be uniquely constrained by the available observables alone. However, we assessed whether meaningful constraints can still be obtained on the most relevant parameters by marginalising over the other ones. Our primary parameters of interest are the orbital period ($P$) and the unseen companion mass ($M_{\text{CO}}$). The remaining orbital parameters – inclination ($i$), eccentricity ($e$), longitude of the ascending node ($\Omega$), argument of periastron ($\omega$), time of periastron passage ($T_{\text{p}}$), and parallax ($\varpi$) – were treated as nuisance parameters required for forward modelling but marginalised in the final inference. Although parallaxes are provided in the Gaia DR3 catalogue, we treated $\varpi$ as a free parameter. This allowed parallax uncertainties to be consistently propagated into the inferred orbital parameters and companion masses, which is particularly important for systems with low parallax significance.
@@
where ‘obs’ denotes the catalogued summary statistics and ‘pred’ the model predictions for a given parameter set $\@vec{\theta}$. The likelihood measures discrepancies in units of scatter, given by $\epsilon_{\texttt{ruwe}}$ and $\epsilon_{\sigma_{\text{RV}}}$, and penalises them in log-space. The use of logarithmic quantities is empirically motivated; forward modelling of NSS systems shows that the combined scatter – from intrinsic variation and model uncertainty – is constant in log-space and can be approximated by a Gaussian distribution (see Fig. 2). Expressing the likelihood in terms of $\ln\mathtt{ruwe}$ and $\ln\sigma_{\mathrm{RV}}$ also ensures a symmetric treatment of relative over- and under-predictions.
@@
We adopted log-uniform priors for period and companion mass, and uniform priors for all other orbital parameters over physically motivated ranges: $\log P/\mathrm{d}\in[1,4]$, $\log M_{\text{CO}}/M_{\odot}\in[\log(0.05),\log(30)]$, $\cos i\in(0,1]$, $e\in[0,0.95]$, $\Omega,\omega\in[0,2\pi)$, and $T_{p}/\text{d}\in[0,P]$. For the parallax, we adopted a normal prior centred on the Gaia DR3 value after applying the global zero-point correction of Lindegren et al. (2021a). The prior standard deviation was set by the reported parallax uncertainty, inflated as a function of ruwe following the prescription of El-Badry (2025), to account for underestimated astrometric errors in unresolved binaries. The flux ratio was fixed to zero, which is appropriate for dark companions. For luminous companions, a non-zero flux ratio would reduce the observed ruwe, leading to underestimated periods and masses. This bias is acceptable since we target dark, massive companions, and for luminous giants, the flux contribution from stellar companions is typically negligible.
@@
We sampled posterior distributions using 32 walkers and 2,500 steps with a 500 step burn-in phase. Convergence assessment is challenging due to the underconstrained problem and high shot noise (e.g. from stochastic Gaia RV observation date sampling). We empirically evaluated convergence on a subset of stars by verifying that: (a) randomly initialised walkers converge to consistent posterior distributions, (b) marginal distributions of $P$ and $M_{\text{CO}}$ remain stable between the second and final thirds of the chain, and (c) increasing chain length by a factor of 10 does not significantly alter results. The inference is computationally efficient, requiring approximately 10 seconds per star on a single CPU core. We report posterior-weighted mean values of $P$ and $M_{\text{CO}}$, marginalised over nuisance parameters.
@@
Figure 3 illustrates the results of this MCMC fitting procedure for two test binary systems: a stellar binary (left panels) and a system containing a compact object (right panels), specifically Gaia BH 2 (El-Badry et al., 2023a). The upper panels display the marginalised posterior probability distribution function (PDF) in the plane of orbital period vs companion mass. The orange star denotes the true input parameters of each system, while the blue star marks the inferred best-fit parameters, defined as the posterior-weighted mean. The dashed ellipses show 1-$\sigma$ and 2-$\sigma$ confidence contours, derived from the posterior covariance matrix. Above each panel, we show the marginalised 1-D posterior PDF of the companion mass, normalised so that the sum of all posterior weights is unity. The bottom panels show the predicted observables ruwe and $\sigma_{\text{RV}}$ for each MCMC sample, colour coded by the corresponding companion mass. The orange star shows the observed values (from the Gaia source catalogue).
@@
The first system (Gaia DR3 4609711364265062016) appears in the NSS catalogue as a AstroSpectroSB1 binary with an orbital period of 141 d. The Gaia binary mass catalogue assigns it a primary mass of $\sim 1.1\,M_{\odot}$ and a companion mass of $\sim 0.35\,M_{\odot}$. The observed values of ruwe and RV uncertainty are modest ($\texttt{ruwe}=1.81$, $\sigma_{\text{RV}}=2.10$ km s-1), consistent with a luminous stellar companion. Our inference correctly recovers a low-mass companion in a short-period orbit ($P_{\text{fit}}=120^{+80}_{-50}\,\text{d},\,\,M_{\text{CO, fit}}=0.4\pm 0.2\,\text{M}_{\odot}$). Small mismatches between the true and inferred values and large uncertainties are expected, given the degeneracies in orbital orientation and eccentricity.
@@
The second example is Gaia BH 2 (Gaia DR3 5870569352746779008), a giant star orbited by a $\sim 9\,\text{M}_{\odot}$ BH in a long-period eccentric orbit ($P\approx 1277$ d, $e\approx 0.53$; El-Badry et al., 2023a). It has a high ruwe value of $9.22$, with moderate RV uncertainty ($\sigma_{\text{RV}}=2.08$ km s-1). Matching both observables requires a high companion mass, and our inference yields a posterior peak at $P_{\text{fit}}=1200^{+1000}_{-600}\,\text{d}$ and $M_{\text{CO, fit}}=7^{+8}_{-4}\,\text{M}_{\odot}$, consistent with the true values and a dark remnant companion. For systems with orbital periods longer than the DR3 baseline, there is a degeneracy between period and companion mass; a longer-period, more massive companion can produce similar observables (ruwe, $\sigma_{\text{RV}}$) as a shorter-period, less massive one due to limited orbital coverage. We capped the prior at $30\,\text{M}_{\odot}$, which constrained the posterior to a physically motivated BH mass regime, but also truncated the distribution. For candidates like BH 2 that lie near this boundary, the inferred companion masses should therefore be interpreted as lower limits under the assumed log-uniform prior.
@@ 4 A catalogue of RGB+BH candidates
We are now in a position to apply this forward-modelling technique to a sample of evolved stars from Gaia DR3. Giant stars offer several advantages in searches for compact companions: (1) with their high luminosities they are likely to outshine their companions, resulting in flux ratios near zero and photocentre motions dominated by the primary; and (2) while stars low on the RGB may still host companions in close ($\sim$ weeks) orbits, the large current or past radii of more evolved giants preclude short-period orbits, reducing contamination from hierarchical triples, a major source of false positives in main-sequence binary searches (Andrew et al., 2022). A corresponding main-sequence sample is provided in Appendix B.
@@ 4.1 Sample selection and filtering
These cuts reduced the sample to 37,421 stars. The RV amplitude threshold roughly corresponds to $\sigma_{\mathrm{RV}}\gtrsim 2$ km s-1. These thresholds are motivated by two considerations: (a) In broader tests, systems with low RV scatter and low ruwe – which occupy the lower left region of the bottom panels in Fig. 3 – were almost exclusively binaries with low-mass companions. Due to the rarity of BHs and the high contamination in this region, we focussed on areas of parameter space where BHs exhibit stronger and more distinct dynamical signatures. (b) Although BHs in wide or low-inclination orbits may be excluded by these cuts, we found that the impact on overall completeness is small for the period range of interest (see Sect. 4.6).
To further refine the sample, we cross-matched with the stellar parameter catalogue from Zhang et al. (2023), which provides $T_{\mathrm{eff}}$, $\log g$, and [Fe/H] estimates from BP and RP (XP) spectra for 220 million stars. We retained stars with $T_{\mathrm{eff}}<6000$ K to remove reddened main-sequence stars, and required reliable parameter estimates (quality_flags ¡ 8). This removed very red supergiants and reduced the sample to 21,028 evolved stars.
@@
Figure 4: CMD of the RGB candidate sample (2D histogram), overlaid on all stars within 100 pc from Gaia DR3 (grey points). Highlighted are the confirmed BH binaries Gaia BH 2 and BH 3, and the candidate LAMOST BH. Neutron star binaries (triangles) and Gaia BH 1 are shown for reference, but lie outside our RGB selection.